find the y-intercept of a line that passes through (4,-6) and has a slope of -1.
Answer by mananth(16946)

m= -1

Plug value of the slope and point ( 4 , -6 ) in
Y = m x + b
-6.00 = -4 + b
b= -6 - -4
b= -2
So the equation will be
Y = -x -2
the y intercept =-2
m.ananth@hotmail.ca